# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v2

@@ Abstract
Diffusion Models achieve state-of-the-art performance in generating new samples but lack low-dimensional latent space that encodes the data into meaningful features. Inversion-based techniques try to solve this issue by reversing the denoising process and mapping images back to their approximated starting noise. In this work, we thoroughly analyze this procedure and focus on the relation between the initial Gaussian noise, the generated samples, and their corresponding latent encodings obtained through the DDIM inversion. First, we show that latents exhibit structural patterns in the form of less diverse noise predicted for smooth image regions. Next, we explain the origin of this phenomenon, demonstrating that, during the first inversion steps, the noise prediction error is much more significant for the plain areas than for the rest of the image. Finally, we present the consequences of the divergence between latents and noises by showing that the space of image inversions is notably less manipulative than the original Gaussian noise. This leads to a low diversity of generated interpolations or editions based on the DDIM inversion procedure and ill-defined latent-to-image mapping. Code is available at https://github.com/luk-st/taba.
@@ 1 Introduction
Diffusion-based probabilistic models (DMs) [43], have achieved state-of-the-art results in many generative domains including image [7], speech [37], video [15], and music [28] synthesis. Nevertheless, one of the significant drawbacks that distinguish diffusion-based approaches from other generative models like Variational Autoencoders (VAEs) [23] is the lack of implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations. To mitigate this issue, several works seek meaningful relations in the starting noise used for generations. This method, known as an inversion technique, was introduced with Denoising Diffusion Implicit Models (DDIM) [44], but led to the proliferation of works [10, 32, 17, 29, 16, 35]. The core idea is to use a denoising diffusion model to predict noise and add it to the image instead of subtracting it. Repeating this process multiple times effectively traces the backward diffusion trajectory, approximating the original noise that could have generated the image. However, due to approximation errors and biases introduced by the trained denoising model, discrepancies arise between the initial Gaussian noise and the reconstructed latent representation.
@@
Figure 1: DDIM inversion produces latent encodings that exhibit less diverse noise in the smooth image areas than in the non-plain one. Through in-depth analysis, we attribute this problem to the errors of noise prediction in the first inversion steps.
While recent works [10, 32, 35, 17, 58] try to improve the inversion procedure from the perspective of tasks such as image reconstruction, editing or interpolation, in this work, we propose to focus solely on the inversion mechanism itself and closely study the errors that DDIM inversion introduces.
@@
To that end, we analyze the relation between sampled Gaussian noise, generated images, and their inverted latent encodings. First, we review existing studies and conduct additional analyses demonstrating that the reverse DDIM technique produces latent representations with pixel correlations that deviate significantly from a Normal distribution. We then experimentally show that this deviation manifests as lower diversity in latent codes, particularly in regions corresponding to smooth image surfaces. To further investigate the cause of this phenomenon, we find that the approximation error of DDIM inversion is higher for pixels representing smooth areas than for other parts of the image. Additionally, we show that this error discrepancy occurs predominantly during the initial steps of the inversion process.
To present the consequences of the observed divergence, we analyze the effects of using latent embeddings instead of raw noise in practical DM applications. Our results show that the latent space is less adaptable for manipulation compared to the noise space. This limitation is particularly noticeable in lower-quality interpolations between randomly chosen latents and less expressive edits, especially in smooth regions of source images. In our final experiment, we revisit the previous observation to highlight that this issue primarily arises during the early stages of the image inversion process, as replacing just $6\%$ of initial steps with ground truth values from the backward diffusion pass significantly mitigates observed limitations.
Our main contributions can be summarized as follows:
@@
  • We show that reverse DDIM produces latent representations that deviate from Normal distribution mostly because of less diverse values for the parts of the latents corresponding to the plain image surfaces.
@@
  • We explain the origins of this phenomenon, showing that the inversion approximation error is more notable for the plain pixel areas, especially during the first noising steps.
@@ 2 Background and Related Work
The training of DMs consists of forward and backward diffusion processes, where in the context of Denoising Diffusion Probabilistic Models (DDPMs [14]), the former one with training image $x_{0}$ and a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be expressed as
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (1)
with $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$.
@@
$\begin{split}x_{t-1}&=\sqrt{\bar{\alpha}_{t-1}}\cdot\underbrace{(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon^{(t)}_{\theta}(x_{t},c))/\sqrt{\bar{\alpha}_{t}}}_{x_{0}\text{ prediction}}\\ &+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(x_{t},c)}_{\text{direction pointing to }x_{t}}+{\sigma_{t}z_{t}},\end{split}$ (2)
where $\epsilon^{(t)}_{\theta}(x_{t},c)$ is an output of a neural network (such as U-Net), and can be expressed as a combination of clean image ($x_{0}$) prediction, a direction pointing to previous denoising step ($x_{t}$), and a stochastic factor $\sigma_{t}z_{t}$, where $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$.
In the standard DDPM model [14], the $\eta$ parameter is set to $\eta=1$. However, changing it to $\eta=0$ makes the whole process a deterministic Denoising Diffusion Implicit Model (DDIM) [44], a class of DMs we target in this work.
@@
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal and Nichol [7] propose to approximate this equation by assuming the local linearity between directions $x_{t-1}$ to $x_{t}$ and $x_{t}$ to $x_{t+1}$, which implies that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance [13, 32] for better prompt adherence. As a result, also noticed by recent works [10, 35], latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing pixel correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables:
  • Gaussian noise variable, ${\mathbf{x_{T}}}$, used as an input to generate an image through a backward diffusion process.
  • Image sample, ${\mathbf{x_{0}}}$, the outcome of the generation process produced by the Diffusion Model.
  • Latent encoding, ${\mathbf{\hat{x}_{T}}}$, the result of the DDIM inversion procedure as introduced in Eq. 3.
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting [53], interpolation [7, 58], and edition [47, 21, 12, 3, 6].
Several works [32, 10, 17, 29, 16, 31, 11, 4, 8, 57, 35, 48, 49, 34, 50, 2, 27, 18] aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) [13]. Null-text inversion [32] extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques, improve inversion for image editing through seeking embeddings [31, 11, 8] or leveraging DDIM latents [4] for guidance. On the other hand, some works leverage additional numerical methods [29, 34, 10, 16] to minimize inversion error. In particular, Renoise [10] iteratively improves the prediction of added noise in each diffusion step via predictor-corrector technique. Huberman-Spiegelglas et al. [17] followed by Brack et al. [2] propose inversion methods for non-probabilistic diffusion models (DPMs), enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. [35] propose to additionally regularize final DDIM Inversion outputs for better image editing, while Lin et al. [27] introduce alternative noise scheduler to improve inversion stability.
Noise-to-image mapping in diffusion models.
While in this work, we focus mostly on the differences between random noise, associated generations, and their inversions, several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. [19] show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. [54], observing that even models with different architectures converge to the same score function and, hence, the same noise-to-generations mapping. Khrulkov and Oseledets [20] show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works [22, 25] contradict this result. In this work, we extend the analysis further by cross-checking mappings between images with their noises or latents created with DDIM inversion.
@@ 3 Experiments
In this section, we provide a series of empirical observations on the relation between sampled noises, their generations, and latent representations.
3.1 Experimental details
We employ diffusion models working directly in the pixel space or in the latent space of the autoencoder. In the case of the former ones, we use UNet-based architectures, including three unconditional Ablated Diffusion Models (ADMs, [7]) in three different image resolution settings: $32\times 32$, $64\times 64$, and $256\times 256$, and Deepfloyd IF [46], text-to-image generative model sampling images in $64\times 64$ resolution. Additionally, we explore the unconditional Latent Diffusion Model (LDM, [39]) with a U-Net model as a backbone and class-conditional Diffusion Transformer (DiT, [36]), which performs diffusion process in the latent space as well. Studied models were trained on diverse datasets, including CIFAR-10, ImageNet, CelebA-HQ, and LAION-A, as presented in Tab. 1. We use original trained checkpoints provided by the authors of those works. Both DiT and IF are conditional models that employ CFG for better prompt adherence. For conditioning, we provide prompts from Recap-COCO-30K dataset [26] for IF and ImageNet class names for DiT. However, as noted by Mokady et al. [32], cfg introduces additional error to the DDIM inversion, so to focus solely on the inversion approximation error in our setup, we disable cfg by setting the guidance scale to $w=1$.
Property ADM-32 ADM-64 ADM-256 LDM DiT IF
@@ 3.1 Experimental details
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in both pixel and latent spaces.
For sampling, we use the DDIM sampler and, unless stated otherwise, set $T=100$ diffusion steps for both sampling and inverting to the latent encodings.
@@
Reversed sampling (inversion) process provides the foundation for methods created for diverse applications. The underlying assumption of those methods is that by encoding the image back with a denoising model, we can obtain the original noise that can be used to reconstruct or manipulate the image. In this work, we challenge this assumption and ask the first question:
@@
Prior work and findings.
Several recent works highlight that outputs of the DDIM inversion differ from the standard Gaussian noise [35, 10] and that the difference can be attributed to the approximation error [16, 49, 27]. In particular, Parmar et al. [35], followed by Garibi et al. [10], observe that latents “do not follow the statistical properties of uncorrelated, Gaussian white noise, causing poor editability”.
Experiments.
First, we confirm existing claims on the presence of correlations in the latent encodings (${\mathbf{\hat{x}_{T}}}$) by comparing them to images (${\mathbf{x_{0}}}$) and noises (${\mathbf{x_{T}}}$) across diverse diffusion architectures. In Tab. 2, we calculate a mean of top-$20$ Pearson correlation coefficients (their absolute values) inside $C\times 8\times 8$ pixel patches, where $C$ is a number of channels (pixel RGB colors or latent space dimensions for latent models). The results, calculated with $10000$ samples, confirm that latent representations have significantly more correlated pixels than the independent Gaussian noises.
@@
Table 2: Mean of top-20 Pearson correlation coefficients inside $8\times 8$ patches for random Gaussian noises, latent encodings, and generated samples. Latents created with the reverse DDIM procedure are much more correlated than noises.
In Fig. 2, we show how the measured correlations manifest themselves in the latent codes calculated with inverse DDIM. For pixel models such as Deepfloyd IF, we observe clear groups of correlated pixels as presented in Fig. 2. For more complex diffusion architectures, such as latent diffusion models, we can highlight the inversion error by plotting the difference between the latent and the noise, as presented in Fig. 2. This property holds also for LDMs with a $4$-channel latent space, where Principal Component Analysis (PCA) is used to present the differences between the latents and noise for the DiT model (Fig. 2).
@@
Figure 2: Latent encodings exhibit image patterns. For small pixel-space models (a), we can observe correlations directly in the inversion results. For bigger models, especially LDMs, the same patterns can be observed in the absolute errors between the latents and the noise (b). This observation also holds for LDM models operating on $4$-channels, where we use PCA for visualization (c).

### v3

@@ Abstract
Diffusion Models achieve state-of-the-art performance in generating new samples but lack a low-dimensional latent space that encodes the data into meaningful features. Inversion-based methods address this by reversing the denoising trajectory, mapping each image back to its approximated starting noise. In this work, we thoroughly analyze this procedure and focus on the relation between the initial Gaussian noise, the generated samples, and their corresponding latent encodings obtained through the DDIM inversion. First, we show that latents exhibit structural patterns in the form of less diverse noise predicted for smooth image regions. As a consequence of this divergence, we present that the space of image inversions is notably less manipulative than the original Gaussian noise. Next, we explain the origin of the phenomenon, demonstrating that, during the first inversion steps, the noise prediction error is much more significant for the plain areas than for the rest of the image. As a surprisingly simple solution, we propose to replace the first DDIM Inversion steps with a forward diffusion process, which successfully decorrelates latent encodings, leading to higher quality editions and interpolations. The code is available at https://github.com/luk-st/taba.
@@ 1 Introduction
Figure 1: DDIM inversion produces latent encodings that exhibit less diverse noise in the smooth image areas than in the non-plain one. We attribute this problem to the errors of noise prediction in the first inversion steps.
Diffusion-based probabilistic models (DMs) [49], have achieved state-of-the-art results in many generative domains including image [8], speech [43], video [17], and music [34] synthesis. Nevertheless, one of the significant drawbacks that distinguishes diffusion-based approaches from other generative models like Variational Autoencoders (VAEs) [26] is the lack of an implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations. To mitigate this issue, several works seek meaningful relations in the starting noise used for generations. This method, known as an inversion technique, was introduced with Denoising Diffusion Implicit Models (DDIM) [50], but led to the proliferation of works [12, 38, 19, 35, 18, 41]. The core idea is to use a denoising diffusion model to predict noise and add it to the image instead of subtracting it. Repeating this process effectively traces the backward diffusion trajectory, approximating the original noise that could have generated the image. However, due to approximation errors and biases introduced by the trained denoising model, discrepancies arise between the initial Gaussian noise and the reconstructed latent representation.
@@
While recent works [12, 38, 41, 19, 63] try to improve the inversion procedure from the perspective of tasks such as image reconstruction, editing, or interpolation, in this work, we focus on the inversion process itself and analyze the errors DDIM inversion introduces. To that end, we analyze the relation between sampled Gaussian noise, generated images, and their inverted latent encodings. First, we review existing studies and conduct additional analyses demonstrating that the reverse DDIM technique produces latent representations with pixel correlations that deviate significantly from a Normal distribution. As presented in fig. 1, we experimentally show that this deviation manifests as lower diversity in latent codes, particularly in regions corresponding to smooth image surfaces. We further attribute this discrepancy to the observation that the approximation error of DDIM inversion is higher for pixels representing smooth areas than for other parts of the image. Additionally, we show that this error discrepancy occurs predominantly during the initial steps of the inversion process. To present the consequences of the observed divergence, we analyze the effects of using latent variables instead of raw noise in practical DM applications. Our results show that the latent space is less adaptable for manipulation than the noise space. This limitation is particularly noticeable in lower-quality interpolations between randomly chosen latents and less expressive edits, especially in smooth regions of source images.
Finally, to mitigate the issue, we propose a surprisingly simple solution, which is replacing the first steps of the DDIM inversion process with a forward diffusion. In the final experiments, we show that such an approach successfully decorrelates the resulting latents, mitigating observed limitations, without degrading the reconstruction quality. Our main contributions can be summarized as follows:
@@
  • We show that reverse DDIM produces latent representations that deviate from the Normal distribution mostly because of less diverse values for the parts of the latents corresponding to the plain image surfaces.
@@
  • We explain the origin of this phenomenon, showing that the inversion approximation error is more notable for the plain pixel areas, especially during the first noising steps.
@@
  • We propose a simple yet effective solution to the problem by substituting early DDIM inversion steps with forward diffusion.
@@ 2 Background and Related Work
The training of DMs consists of forward and backward diffusion processes, where in the context of Denoising Diffusion Probabilistic Models (DDPMs [16]), the former one with training image $x_{0}$ and a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ with $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$.
@@
$\begin{split}x_{t-1}&=\sqrt{\bar{\alpha}_{t-1}}\cdot\underbrace{(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon^{(t)}_{\theta}(x_{t},c))/\sqrt{\bar{\alpha}_{t}}}_{x_{0}\text{ prediction}}+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(x_{t},c)}_{\text{direction pointing to }x_{t}}+{\sigma_{t}z_{t}},\end{split}$ (1)
where $\epsilon^{(t)}_{\theta}(x_{t},c)$ is an output of a neural network (such as U-Net), and can be expressed as a combination of clean image ($x_{0}$) prediction, a direction pointing to previous denoising step ($x_{t}$), and a stochastic factor $\sigma_{t}z_{t}$, where $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$. In the standard DDPM model [16], the $\eta$ parameter is set to $\eta=1$. However, changing it to $\eta=0$ makes the whole process a deterministic Denoising Diffusion Implicit Model (DDIM) [50], a class of DMs we target in this work.
@@
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal and Nichol [8] propose to approximate this equation by assuming the local linearity between directions $(x_{t-1}\rightarrow x_{t})$ and $(x_{t}\rightarrow x_{t+1})$, so that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance [15, 38] for better prompt adherence. As a result, also noticed by recent works [12, 41], latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables: Gaussian noise (${\mathbf{x_{T}}}$, an input to generate an image through a backward diffusion process), image sample (${\mathbf{x_{0}}}$, the outcome of the diffusion model generation process), and latent encoding (${\mathbf{\hat{x}_{T}}}$, the result of the DDIM inversion procedure as introduced in eq. 2).
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting [58], interpolation [8, 63], and edition [53, 23, 14, 3, 7]. Several works [38, 12, 19, 35, 18, 37, 13, 4, 9, 62, 41, 54, 55, 40, 56, 2, 32, 20] aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) [15]. Null-text inversion [38] extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques improve inversion for image editing through seeking embeddings [37, 13, 9] or leveraging DDIM latents [4] for guidance. On the other hand, some works leverage additional numerical methods [35, 40, 12] to minimize inversion error. In particular, Renoise [12] iteratively improves the estimation of the next point along the forward diffusion trajectory by averaging multiple noise predictions with an additional patch-level regularization term penalizing pixel pairs’ correlations to ensure the latents’ editability. Huberman-Spiegelglas et al. [19] followed by Brack et al. [2] propose inversion methods for DDPMs, enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. [41] propose to additionally regularize final DDIM Inversion outputs for better image editing, Lin et al. [32] introduce an alternative noise scheduler to improve inversion stability, while Hong et al. [18] propose an exact inversion procedure for higher-order DPM-Solvers, solving the optimization problem at each step.
@@ 3 Experiments
ADM-32 ADM-64 ADM-256 LDM DiT IF
@@
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in pixel and latent spaces.
For thorough evaluation, we employ six different diffusion models working directly in the pixel space or latent space of the autoencoder described in table 1. For conditioning, we take prompts from the Recap-COCO-30K dataset [31] for IF and ImageNet class names for DiT. However, as noted by Mokady et al. [38], cfg introduces additional errors to the DDIM inversion, so to focus solely on the inversion approximation error in our setup, we disable cfg by setting the guidance scale to $w=1$. For both generation and inversion processes, we use the DDIM sampler with, unless stated otherwise, $T=100$ sampling steps. This choice [^1] We provide more discussion on the impact of sampling steps on observed inversion errors in appendix E. aligns with prior works [18, 12, 24] using from $50$ up to $200$ inversion steps as a proper balance between reconstruction quality and short algorithm runtime.
@@
Prior work and findings.
The reversed sampling (inversion) process provides the foundation for practical methods in many applications. The underlying assumption of those methods is that by encoding the image back with a denoising model, we can obtain the original noise that can be used for reconstruction. However, this assumption is not always fulfilled, which leads to the first question:
@@
Table 2: Mean of top-20 Pearson correlation coefficients inside $8\times 8$ patches for random Gaussian noises, latent encodings, and generations. Latents created with the reverse DDIM are much more correlated than noises.
@@
Figure 2: Latent encodings exhibit image patterns. For small pixel-space models (a), we observe correlations directly in the inversion results. For larger models (e.g., LDMs), the same patterns can be observed in the absolute errors between the latent and noise (b). This observation also holds for LDM models operating on $4$-channels, where we use PCA for visualization (c).
This question relates to several observations from the existing literature, which highlight that outputs of the DDIM inversion differ from the standard Gaussian noise [41, 12] and that the difference can be attributed to the approximation error [18, 55, 32]. In particular, Parmar et al. [41], followed by Garibi et al. [12], observe that latents “do not follow the statistical properties of uncorrelated, Gaussian white noise, causing poor editability”.
Experiments.
First, we consolidate existing claims on the presence of correlations in the latent encodings (${\mathbf{\hat{x}_{T}}}$) by running an initial experiment that compares latents to images (${\mathbf{x_{0}}}$) and noises (${\mathbf{x_{T}}}$) across diverse diffusion architectures. In table 2, we calculate a mean of top-$20$ Pearson correlation coefficients (their absolute values) inside $C\times 8\times 8$ pixel patches, where $C$ is the number of channels (pixel RGB colors or latent space dimensions for latent models). The results, calculated with $10000$ samples, confirm that latent representations have significantly more correlated pixels than the independent Gaussian noises. In fig. 2, we show how the measured correlations manifest themselves in the latent codes calculated with inverse DDIM. For pixel models such as Deepfloyd IF, we observe clear groups of correlated pixels as presented in fig. 2. For latent diffusion models, we can highlight the inversion error by plotting the difference between the latent and the noise, as presented in fig. 2. This property also holds for LDMs with a $4$-channel latent space, where we use PCA to show the differences between the latents and noise for the DiT model (fig. 2).